# A Sampling-based Framework for Hypothesis Testing on Large Attributed Graphs [Scalable Data Science]

### v1

@@ 4.3. PHASE vs $\text{PHASE}_{\text{opt}}$
We evaluate the effectiveness of optimizations in $\text{PHASE}_{\text{opt}}$ compared to PHASE using the DBLP dataset, as shown in Table 2. $\text{PHASE}_{\text{opt}}$ is at least 20 times faster than PHASE while maintaining comparable accuracy (less than 5% accuracy reduction). Consequently, we exclusively use $\text{PHASE}_{\text{opt}}$ for subsequent experiments.
@@ 4.4. Significance
To examine statistical significance and estimation precision, we analyze the trend of p-values and confidence intervals (CI) with increasing $B$. We illustrate this trend using a path hypothesis on DBLP as all other hypotheses show similar trends. As shown in Figure 3(a), the p-value of $\text{PHASE}_{\text{opt}}$ consistently stays below the significance level (e.g., 0.05) and decreases gradually. This trend indicates stronger evidence against the null hypothesis, showing higher statistical significance as $B$ increases.
Figure 3(b) shows narrowing CI with increasing $B$, indicating more precise and reliable estimations of the hypothesis estimator. $\text{PHASE}_{\text{opt}}$ consistently achieves the narrowest bound at any $B$, indicating the most precise and reliable estimate among all samplers.
@@
Figure 5. Accuracy versus sampling proportion plots for DBLP path hypotheses with length three (a) and four (b).
@@ 4.5. Accuracy
Table 3 presents the accuracy performance of eleven existing samplers and $\text{PHASE}_{\text{opt}}$ at a specific sampling proportion, shown in column three. We select this proportion to balance stability and discriminative power, ensuring the accuracy is sufficiently stabilized without saturating across all samplers. For each dataset and type of hypothesis, the accuracy is the average accuracy of three hypothesis examples, each derived from an average of 30 runs. The highest accuracy in each row is highlighted in bold.
@@
In Table 3, $\text{PHASE}_{\text{opt}}$ demonstrates robust performance in all cases, except for MovieLens edge hypotheses, in where $\text{PHASE}_{\text{opt}}$ slightly trails behind NBRW. Comparing among columns, some random walk based samplers, namely SRW, RWR, and ShortestPathS, exhibit good accuracy performance for edge and path hypothesis. On the contrary, RES, RNS, and DBS show ineffectiveness in handling path hypotheses. These observations align with their sampling mechanisms. These node and edge samplers can hardly preserve the path information from $\mathcal{G}$ in the sampled graph. Additionally, comparing $\text{PHASE}_{\text{opt}}$ and FrontierS, $\text{PHASE}_{\text{opt}}$’s superior performance indicates the efficacy of the two weight functions in enhancing the sampler’s hypothesis-awareness to sample relevant nodes, edges, and paths.
(a) Time $(l=3)$
(b) Time ($l=4$)
Figure 6. Time versus accuracy plots for DBLP path hypotheses with length three (a) and four (b).
(a) DB-N1
(b) DB-N2
(c) DB-N3
(d) DB-N3
(e) DB-E1
(f) DB-E2
(g) DB-E3
(h) DB-E3
(i) DB-P1
(j) DB-P2
(k) DB-P3
(l) DB-P3
Figure 7. Comparison of the top five sampling methods for accuracy and time (d, h, l) efficiency across three node (a, b, c), three edge (e, f, g), and three path hypotheses (i, j, k) for the DBLP dataset.
@@
We establish an accuracy ranking for all samplers by averaging their accuracy performance per column in Table 3 and select the top five: $\text{PHASE}_{\text{opt}}$, NBRW, ShortestPathS, RWR, and SRW. These samplers’ accuracy convergence on individual hypotheses is compared in Figures 7 and 8 for DBLP and Yelp datasets, respectively. The figure for MovieLens is excluded due to space constraints and the similarity of results. In these figures, a-c, e-g, and i-k display the accuracy performance of three node, three edge, and three path hypotheses, respectively. Hypotheses of the same type are structured to have progressively fewer relevant nodes, edges, or paths in $\mathcal{G}$. The x-axis is truncated to clearly show the convergence.
In Figures 7 and 8, $\text{PHASE}_{\text{opt}}$ consistently outperforms other methods across all sampling proportions for most hypotheses. For hypotheses of the same type, $\text{PHASE}_{\text{opt}}$, with hypothesis-awareness, consistently demonstrates robust accuracy performance for the hypotheses with fewer relevant nodes, edges, or paths in $\mathcal{G}$. This is particularly evident for DBLP edge and path hypotheses. When relevant nodes, edges, and paths in $\mathcal{G}$ are abundant, conventional sampling methods may have sampled a sufficient number of relevant nodes, edges, or paths to achieve high accuracy, which can diminish the distinctive advantage of $\text{PHASE}_{\text{opt}}$.
Delving into Figure 7, $\text{PHASE}_{\text{opt}}$ dominates in node and path hypotheses for the DBLP dataset. However, ShortestPathS slightly outperforms $\text{PHASE}_{\text{opt}}$ in edge hypotheses one and two, possibly due to the high betweenness centrality of these edges, implying that many shortest paths traverse those edges. Moving to Figure 8, $\text{PHASE}_{\text{opt}}$ demonstrates clear superiority across all types of hypotheses for the Yelp dataset. Nonetheless, ShortestPathS occasionally exhibits competitive performance compared to $\text{PHASE}_{\text{opt}}$ in path hypotheses, which may also be attributed to the betweenness centrality of those paths in $\mathcal{G}$.
Figure 5 shows how path length affects accuracy using the DBLP dataset. $\text{PHASE}_{\text{opt}}$ enables faster convergence of accuracy with longer path lengths, outperforming other samplers.
@@ 4.6. Scalability
(a) YP-N1
(b) YP-N2
(c) YP-N3
(d) YP-N3
(e) YP-E1
(f) YP-E2
(g) YP-E3
(h) YP-E3
(i) YP-P1
(j) YP-P2
(k) YP-P3
(l) YP-P3
Figure 8. Comparison of the top five sampling methods for accuracy and time (d, h, l) efficiency across three node (a, b, c), three edge (e, f, g), and three path hypotheses (i, j, k) for the Yelp dataset.
In Table 4, samplers’ execution time increases as the dataset becomes larger. Among them, RNS generally exhibits the shortest execution time as it uniformly samples nodes randomly. The performance of $\text{PHASE}_{\text{opt}}$ varies across datasets. For the Yelp dataset, it ranks among the highest, particularly for the path hypothesis. This can be attributed to the limited node types in Yelp, resulting in longer extraction times for condensed relevant paths from $\mathcal{S}$. For the DBLP dataset, $\text{PHASE}_{\text{opt}}$ ranks among the lowest five. However, instead of considering execution time alone, we should consider both execution time and accuracy performance to evaluate a sampler’s efficiency and effectiveness.
For the top five samplers ranked by their accuracy performance, we plot their accuracy versus execution time in subfigures d, h, and l of Figures 7 and 8. Due to space limitations, only the third hypothesis, which consists of the least number of relevant nodes, edges, or paths in the input graph, of each type is included. First of all, we observe that the curve of $\text{PHASE}_{\text{opt}}$ is always at the bottom among others. This means that $\text{PHASE}_{\text{opt}}$ consistently achieves high accuracy in the shortest amount of time across all plots. Additionally, its execution time does not exhibit exponential growth compared to other existing hypothesis-agnostic samplers. This can be attributed to $\text{PHASE}_{\text{opt}}$’s hypothesis-awareness, which allows it to efficiently sample relevant nodes, edges, and paths from the input graph to verify the hypothesis.
The impact of path length on accuracy and execution time is illustrated in Figure 6. As the path length increases, $\text{PHASE}_{\text{opt}}$ consistently achieves the highest accuracy within the shortest execution time.
@@ 5. Related Work
There has been a lot of research interest on hypothesis testing (Hunter et al., 2008; Gibbs et al., 2022; Arias-Castro and Verzelen, 2014; Xia and Li, 2017; Croft et al., 2011; Ghoshdastidar and von Luxburg, 2018; Omidvar-Tehrani et al., 2019), proposing sampling techniques on graphs (Leskovec and Faloutsos, 2006; Lee et al., 2012; Li et al., 2019). In this section we summarize the most representative works.
@@
Our goal is more related to works when the object of interest is nodes. However, there are two differences between our work and the existing ones. First, we enable more expressive hypotheses, including edge and path hypothesis, on graphs. Second, we leverage graph sampling methods on graph hypothesis testing.
Sampling Methods on Graphs. Most sampling methods are designed to sample a representative subgraph to accurately estimate graph properties (Leskovec and Faloutsos, 2006; Stutzbach et al., 2009; Maiya and Berger-Wolf, 2010; Hübler et al., 2008). Leskovec and Faloutsos (Leskovec and Faloutsos, 2006) are the first to study this problem in real-world networks (Leskovec and Faloutsos, 2006) by proposing sampling techniques to maintain degree distribution, clustering coefficient and distribution of component sizes. Hübler et al. (Hübler et al., 2008) propose the Metropolis-Hastings sampling method. Later, Maiya and Berger-Wolf (Maiya and Berger-Wolf, 2010) propose a community representativeness sample and design a community structure expansion sampler. Many sampling methods are proposed to improve the efficiency and convergence rate of simple random walk (Lee et al., 2012; Li et al., 2019). Besides representative sampling, other sampling methods are designed for specific tasks, such as graph compression (Besta et al., 2019), community detection(Maiya and Berger-Wolf, 2010), and graph visualization (Rafiei and Curial, 2005; Wu et al., 2017). However, none of them is designed for hypothesis testing on attributed graphs.
@@ 6. Conclusion
In this paper, we developed a framework for hypothesis testing on large attributed graphs. Our framework accommodates eleven common hypothesis-agnostic samplers as well as new hypothesis-aware samplers. We proposed dedicated optimizations to speed up sampling while achieving high test significance and accuracy. We demonstrated theoretically and empirically that the convergence rate of our hypothesis-aware sampling.

### v2

@@ 4.3. PHASE vs $\text{PHASE}_{\text{opt}}$
We evaluate $\text{PHASE}_{\text{opt}}$’s optimizations compared to PHASE in Table 3 and 4. For DBLP, $\text{PHASE}_{\text{opt}}$ is at least 43 times faster than PHASE with minimal accuracy loss (under 4%). Consequently, we use $\text{PHASE}_{\text{opt}}$ exclusively in further experiments and discussions.
@@ 4.4. Significance
To assess statistical significance and estimation precision, we evaluate p-value trends and confidence intervals (CIs) as $B$ increases using DB-P3, as shown in Figure  5. Similar trends are observed for other hypotheses. $\text{PHASE}_{\text{opt}}$ consistently maintains p-values below the significance level (e.g., 0.05), showing stronger evidence against the null hypothesis as $B$ increases. It also exhibits the narrowest CI at all $B$, reflecting the highest precision among other samplers.
@@
(c) Time-Accuracy $(l=3)$
(d) Time-Accuracy $(l=4)$
Figure 6. Accuracy (a, b) and time-accuracy (c, d) performance for DBLP path hypotheses.
(a) DB-N1 (easy)
(b) DB-N2 (medium)
(c) DB-N3 (hard)
(d) DB-N3 (hard)
(e) DB-E1 (easy)
(f) DB-E2 (medium)
(g) DB-E3 (hard)
(h) DB-E3 (hard)
(i) DB-P1 (easy)
(j) DB-P2 (medium)
(k) DB-P3 (hard)
(l) DB-P3 (hard)
Figure 7. Comparison of the top five sampling methods for accuracy (a-c, e-g, i-k) and time-accuracy (d, h, l) performance across three node (a-d), three edge (e-h), and three path hypotheses (i-l) for DBLP.
@@ 4.5. Accuracy
Table 3 presents the accuracy of 11 existing samplers and $\text{PHASE}_{\text{opt}}$ with a fix $B$ in column three. $B$ is set to ensure the accuracy is sufficiently stabilized without saturating across all samplers. Each row reports an average accuracy from three examples, based on 30 runs each. The highest and second highest accuracies are highlighted in bold and underlined, respectively.
@@
Table 3 shows $\text{PHASE}_{\text{opt}}$’s robust performance across most scenarios, except for MovieLens edge hypotheses where it slightly lags behind NBRW. SRW, RWR, and ShortestPathS perform well on edge and path hypotheses, whereas RES, RNS, and DBS struggle with path hypotheses. This observation aligns with their sampling mechanisms: node and edge samplers can hardly preserve the path information from $\mathcal{G}$. Also, $\text{PHASE}_{\text{opt}}$ outperforms FrontierS, showcasing the effectiveness of the two weight functions in enhancing the sampler’s hypothesis-awareness.
@@
We rank the accuracy of all samplers by averaging their accuracy per column in Table  3 and identify the top five: $\text{PHASE}_{\text{opt}}$, NBRW, ShortestPathS, RWR, and SRW. The accuracy performance on individual hypotheses for DBLP and Yelp is shown in Figures  7 and  8, respectively. Figures for MovieLens are omitted due to the space constraints and result similarity. Subfigures a-c, e-g, and i-k depict accuracy versus $B$ for three node, edge, and path hypotheses, respectively, with truncated x-axes to highlight the convergence.
(a) YP-N1 (easy)
(b) YP-N2 (medium)
(c) YP-N3 (hard)
(d) YP-N3 (hard)
(e) YP-E1 (easy)
(f) YP-E2 (medium)
(g) YP-E3 (hard)
(h) YP-E3 (hard)
(i) YP-P1 (easy)
(j) YP-P2 (medium)
(k) YP-P3 (hard)
(l) YP-P3 (hard)
Figure 8. Comparison of the top five sampling methods for accuracy (a-c, e-g, i-k) and time-accuracy (d, h, l) performance across three node (a-d), three edge (e-h), and three path hypotheses (i-l) for Yelp.
In Figures 7 and  8, $\text{PHASE}_{\text{opt}}$ consistently outperforms other samplers across most hypotheses and sampling proportions. When there are abundant relevant nodes, edges, or paths in $\mathcal{G}$ (e.g, easy and medium hypotheses), hypothesis-agnostic samplers can perform well, reducing $\text{PHASE}_{\text{opt}}$’s relative advantage. However, its superiority is more obvious for more difficult hypotheses (e.g., hard hypotheses). Also, for difficult hypotheses with longer paths in DBLP, $\text{PHASE}_{\text{opt}}$ achieves the highest accuracy at any sampling proportion, as shown in Figures 6(a) and  6(b). Moreover, ShortestPathS sometimes shows competitive performance, possibly due to the high betweenness centrality of the concerned nodes, edges, or paths, which implies many shortest paths traversing them.
@@ 4.6. Scalability
In Table  4, the lowest execution time is in bold. First, the execution time increases with dataset size. Second, RNS generally requires the shortest execution time because it uniformly samples nodes. Our method, $\text{PHASE}_{\text{opt}}$, has varying time performance across datasets. Specifically, it ranks among the highest in execution time for Yelp, especially for path hypotheses. This is attributed to the limited node types that extend path extraction times. While for DBLP, its execution time ranks in the lowest five.
As $\text{PHASE}_{\text{opt}}$ can achieve high accuracy with a small $B$, it is unfair to compare the time efficiency at a fixed $B$. Instead, we plot accuracy versus execution time in subfigures d, h, and l of Figures  7 and  8 for DBLP and Yelp, respectively, for the top five samplers ranked by accuracy. Due to the space constraints, only the most difficult hypothesis of each type is shown. We measure time and accuracy starting from $B=1000$ in increments of 1000 until the accuracy reaches one, or time reaches 50s (DBLP) and 30s (Yelp). We find that $\text{PHASE}_{\text{opt}}$ consistently achieves high accuracy in the shortest time and with the least $B$. Also, the execution time does not grow exponentially. Moreover, when the path length increases, as shown in Figure  6(c) and  6(d), $\text{PHASE}_{\text{opt}}$ achieves the highest accuracy in the shortest time.
@@ 5. Related Work
There has been a lot of research interest on hypothesis testing  (Hunter et al., 2008; Gibbs et al., 2022; Arias-Castro and Verzelen, 2014; Xia and Li, 2017; Croft et al., 2011; Ghoshdastidar and von Luxburg, 2018; Omidvar-Tehrani et al., 2019) and graph sampling techniques (Leskovec and Faloutsos, 2006; Lee et al., 2012; Li et al., 2019). In this section, we summarize the most representative works.
@@
Our goal is closely related to works that focus on nodes as the object of interest. However, there are two major differences between our work and the existing ones. First, we enable more expressive hypotheses, including edge and path hypotheses, on graphs. Second, we leverage graph sampling methods to conduct hypothesis testing.
Sampling Methods on Graphs. Most sampling methods are designed to sample a representative subgraph to accurately estimate graph properties (Leskovec and Faloutsos, 2006; Stutzbach et al., 2009; Maiya and Berger-Wolf, 2010; Hübler et al., 2008). Leskovec and Faloutsos  (Leskovec and Faloutsos, 2006) are the first to study this problem in real-world networks (Leskovec and Faloutsos, 2006) by proposing sampling techniques to maintain degree distribution, clustering coefficient, and distribution of component sizes. Hübler et al. (Hübler et al., 2008) propose the Metropolis-Hastings sampling method. Maiya and Berger-Wolf (Maiya and Berger-Wolf, 2010) define the community representativeness sample and propose a community structure expansion sampler. Later, many sampling methods are proposed to improve the efficiency and convergence rate of simple random walk (Lee et al., 2012; Li et al., 2019). Besides representative sampling, some sampling methods are designed for specific tasks, such as graph compression (Besta et al., 2019), community detection (Maiya and Berger-Wolf, 2010), and graph visualization (Rafiei and Curial, 2005; Wu et al., 2017). However, none of them is designed for hypothesis testing on attributed graphs.
@@ 6. Conclusion
In this paper, we develop a framework for hypothesis testing on large attributed graphs, which accommodates 11 existing hypothesis-agnostic samplers and new hypothesis-aware samplers. We propose dedicated optimizations to speed up sampling while achieving high test significance and accuracy. We also demonstrate theoretically and empirically that our hypothesis-aware sampling achieves earlier convergence of the hypothesis estimator than other methods.
@@
Acknowledgements.
Reynold Cheng, Yun Wang, and Chrysanthi Kosyfaki were supported by the University of Hong Kong (Project 109000579), the HKU Outstanding Research Student Supervisor Award 2022-23, and the HKU Faculty Exchange Award 2024 (Engineering).